OpenStudy (anonymous):

Given the two functions below, what is the x-intercept of g(x)? f(x) = x + 14 g(x) = f(-x)

OpenStudy (anonymous):

x-intercept: when y = 0 g(x) = f(-x) = 14 - x = y when y = 0 = 14 - x x = 14